OpenStudy (secret-ninja):

The graph of f(x) = x^2 is shifted 3 units to the left to obtain the graph of g(x). Which of the following equations best describes g(x)? g(x) = (x + 3)^2 g(x) = x^2 - 3 g(x) = x^2 + 3 g(x) = (x - 3)^2

mathslover (mathslover):

Earlier it was \(f(x) = x^2\) for origin to be the reference point. If we shift the graph 3 units to the left, that is we are shifting the x coordinate to the left with respect to the origin.  Clearly, for g(x), \(x' = x-3\) So, g(x) = (x-3)^2

OpenStudy (secret-ninja):

Thank you ^-^